OpenStudy (anonymous):

find the vertex and axis of symmetry. f(x) = -x^2 +6x+ 6

OpenStudy (anonymous):

A formula for the axis of symmetry can be found by simply averaging the roots from the quadratic formula:\[x=\frac{1}{2}\left[\frac{-b+\sqrt{b^2-4ac}}{2a}+\frac{-b-\sqrt{b^2-4ac}}{2a}\right]=\boxed{-\dfrac{b}{2a}}\]The vertex is then where this axis of symmetry intersects the parabola.

OpenStudy (anonymous):

that's awesome yakeyglee ^^^

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

since you know where the axis of symmetry is, you can use that to find the vertex... HINT: the vertex has the same x coordinate as the line of symmetry.

OpenStudy (anonymous):

For any parabola of the form \(ax^2+bx+c=0 \) the vertex will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \). Equation of line of symmetry will be at \(x=-\frac b {2a} \)